OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (valpey):

Is this how you read this:? \[3\sqrt{2x}+x\sqrt{8x}-5\sqrt{18x}=-10x\sqrt{2x}\]

OpenStudy (valpey):

Let's just say Nakim would have been correct if the problem was: \[3x\sqrt{2x}+x\sqrt{8x}-5x\sqrt{18x}=-10x\sqrt{2x}\]

OpenStudy (valpey):

Do you see why that would be?